# MQFL-FHE: Multimodal Quantum Federated Learning Framework with Fully Homomorphic Encryption

### v4

@@ I Introduction
Federated Learning (FL) has emerged as a powerful paradigm for collaboratively training machine learning models across decentralized data sources [1], allowing multiple clients to enhance model performance while preserving data privacy and complying with regulations such as the general data protection regulation (GDPR). FL alone is insufficient to fully protect data privacy, as the global model and updates shared between clients and the central server remain vulnerable to inference attacks. For instance, membership inference attacks use the model parameters or gradients to deduce whether a specific data sample was part of the training dataset [2]. Similarly, attribute inference attacks attempt to uncover sensitive attributes of individual data samples by exploiting correlations between model updates and data features [3]. Fully homomorphic encryption (FHE) was proposed by Gentry to mitigate vulnerabilities by allowing computations on encrypted data without decryption [4], ensuring that sensitive information remains secure even during model training and aggregation within new gen FL. It works well against gradient leakage attacks, in which hackers try to recover private information by looking at the gradients that are sent back and forth during training [5]. Adding FHE to FL adds a lot of extra work to computers, causing delays and less accurate models during the aggregation phase [6]. Consequently, while FHE protects data privacy, it can hinder the overall performance of FL systems, which is crucial for applications requiring high accuracy or real-time predictions. Despite the growing interest in multimodal FL, which aims to leverage various types of data from multiple clients, there exists a gap in research that addresses the unique challenges posed by FHE in this context [7]. Furthermore, the application of quantum computing (QC) to enhance FL frameworks has been largely unexplored. QC has the potential to generate richer representation within the federated system, which could alleviate performance bottlenecks introduced by FHE [8].
@@ II Related Works
QC in FL: Innan et al. introduced the federated quantum neural network (FedQNN) [10], which integrates quantum machine learning (QML) with QFL principles [11]. The experiments were conducted across genomics and healthcare datasets, laying the groundwork for investigations into quantum-enhanced learning methods in privacy-sensitive domains. This has been further advanced by efforts in QFL, which are creating new avenues in QML. In particular, multiple studies have emphasized using QFL in quantum systems and decentralized data frameworks [12, 13]. Applications in healthcare and other domains, using both quantum hardware and classical simulations, have shown promising results [14, 15, 16, 17, 18, 19]. Despite this progress, challenges related to resource allocation and efficient implementation persist [20, 21, 22], highlighting ongoing efforts to address these issues.
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
In MQFL-FHE, we employ an immediate feature-level fusion by utilizing multi-head attention mechanisms to extract relevant features from diverse input modalities [27]. The client-specific preprocessing function is represented as $\mathcal{D}_{k}\leftarrow\text{PrepareMultimodalDataset}(k)$, ensuring that varying input types are effectively managed and allowing local models to learn from different sources. The local model update for client $k$ can be expressed as:
@@ V-A Ablation Study
  • FL+FHE vs QFL+FHE: To evaluate the impact of each module on the final performance of the global model, we conduct experiments on DNA+MRI multimodal datasets for an ablation study. As Fig. 5 shows, the confusion matrices show that QFL+FHE significantly improves classification accuracy compared to FL+FHE. For example, in the DNA dataset, QFL+FHE achieves better diagonal dominance, especially in class 6 with 0.31 accuracy, compared to 0.34 for FL+FHE. In the MRI dataset, QFL+FHE shows superior performance for crucial classes such as glioma and pituitary, with values of 0.26 and 0.2, respectively, compared to 0.17 and 0.03 for FL+FHE. This highlights that QFL+FHE which is FL+FHE when integrated with QC offers improved classification performance and better separation of challenging classes, particularly for more difficult or underrepresented categories.

### v5

@@ I Introduction
Federated Learning (FL) has emerged as a powerful paradigm for collaboratively training machine learning models across decentralized data sources [1], allowing multiple clients to enhance model performance while preserving data privacy and complying with regulations such as the general data protection regulation (GDPR). FL alone is insufficient to fully protect data privacy, as the global model and updates shared between clients and the central server remain vulnerable to inference attacks. For instance, membership inference attacks use the model parameters or gradients to deduce whether a specific data sample was part of the training dataset [2]. Similarly, attribute inference attacks attempt to uncover sensitive attributes of individual data samples by exploiting correlations between model updates and data features [3]. Fully homomorphic encryption (FHE) was proposed by Gentry to mitigate vulnerabilities by allowing computations on encrypted data without decryption [4], ensuring that sensitive information remains secure even during model training and aggregation within next-generation FL. It works well against gradient leakage attacks, in which hackers try to recover private information by looking at the gradients that are sent back and forth during training [5]. Adding FHE to FL adds a lot of extra work to computers, causing delays and less accurate models during the aggregation phase [6]. Consequently, while FHE protects data privacy, it can hinder the overall performance of FL systems, which is crucial for applications requiring high accuracy or real-time predictions. Despite the growing interest in multimodal FL, which aims to leverage various types of data from multiple clients, there exists a gap in research that addresses the unique challenges posed by FHE in this context [7]. Furthermore, the application of quantum computing (QC) to enhance FL frameworks has been largely unexplored. QC has the potential to generate richer representation within the federated system, which could alleviate performance bottlenecks introduced by FHE [8].
@@ II Related Works
QC in FL: Innan et al. introduced the federated quantum neural network (FedQNN) [10], which integrates quantum machine learning (QML) with QFL principles [11, 12, 13]. The experiments were conducted across genomics and healthcare datasets, laying the groundwork for investigations into quantum-enhanced learning methods in privacy-sensitive domains. This has been further advanced by efforts in QFL, which are creating new avenues in QML. In particular, multiple studies have emphasized using QFL in quantum systems and decentralized data frameworks [14, 15]. Applications in healthcare and other domains, using both quantum hardware and classical simulations, have shown promising results [16, 17, 18, 19, 20, 21, 22, 23]. Despite this progress, challenges related to resource allocation and efficient implementation persist [24, 25, 26], highlighting ongoing efforts to address these issues.
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
In MQFL-FHE, we employ an immediate feature-level fusion by utilizing multi-head attention mechanisms to extract relevant features from diverse input modalities [31, 32]. The client-specific preprocessing function is represented as $\mathcal{D}_{k}\leftarrow\text{PrepareMultimodalDataset}(k)$, ensuring that varying input types are effectively managed and allowing local models to learn from different sources. The local model update for client $k$ can be expressed as:
@@ V-A Ablation Study
  • FL+FHE vs QFL+FHE: To evaluate the impact of each module on the final performance of the global model, we conduct experiments on DNA+MRI multimodal datasets for an ablation study. As Fig. 5 shows, the confusion matrices show that QFL+FHE significantly improves classification accuracy compared to FL+FHE. For example, in the DNA dataset, QFL+FHE achieves better diagonal dominance, especially in class 6 with 0.31 accuracy, compared to 0.34 for FL+FHE. In the MRI dataset, QFL+FHE shows superior performance for crucial classes such as glioma and pituitary, with values of 0.26 and 0.2, respectively, compared to 0.17 and 0.03 for FL+FHE. This highlights that QFL+FHE, which is FL+FHE when integrated with QC, offers improved classification performance and better separation of challenging classes, particularly for more difficult or underrepresented categories.